OpenStudy (anonymous):

what would be 5+3x(6-4x3)

OpenStudy (anonymous):

\[5+3*(6-4*3)=8(6-12)=8(-6)=-48\]

OpenStudy (anonymous):

thanks dude for the medal you are the best

OpenStudy (anonymous):

k